OpenStudy (anonymous):

v = 4i + 2j - 5k ,w = 5i + 5j + 2k ,u = 3i + 2j - 5k Find w (v x u).

OpenStudy (shamim):

is it\[w.(v \times u)\]

OpenStudy (shamim):

u hv to put a dot sign after w

OpenStudy (shamim):

\[v \times u=(4i+2j-5k)(3i+2j-5k)=12(i \times i)+8(i \times j)-20(i \times k)+6(j \times i)+4(j \times j)-10(j \times k)\]

OpenStudy (shamim):

\[-15(k \times i)-10(k \times j)+25(k \times k)\]

OpenStudy (shamim):

u know that\[i \times i=1 \times 1\times \sin0=0\]

OpenStudy (shamim):

in the same way \[j \times j=0 \]

OpenStudy (shamim):

\[k \times k=0\]

OpenStudy (shamim):

so the result will b\[8k+20j-6k-10i-15j+10i=5j+2k\]

OpenStudy (shamim):

if u want then i wanna do the rest of this math